INVESTIGATING THE EXCITED ELECTRONIC STATES OF BaOH VIA LASER SPECTROSCOPY AND AB INITIO CALCULATION: FURTHER EVIDENCE OF PERTURBATION FROM THE Ã^{′2}Δ STATE

\underline{J. D. TANDY}, J.-G. WANG, P. F. BERNATH, Department of Chemistry, University of York, Heslington, York, YO10 5DD, UK; J. LIÉVIN, Service de Chimie Quantique et Photophysique, Université Libre de Bruxelles, CPi 160/09, 50 av F.D. Roosevelt, B-1050 Bruxelles, Belgium.

Two bands for the Ã^{′2}Δ − X̃^{2}Σ^{+} transitions of BaOH and BaOD have been rotationally analyzed using high-resolution V-type optical-optical double resonance spectroscopy. BaOH and BaOD molecules were synthesized in a Broida-type oven, using a single mode Ti:Sapphire laser and a single mode dye laser for molecular excitation. The observed spectra mimic a typical ^{2}Π−^{2} Σ^{+} transition, believed to emanate from single or triple quanta of the bending vibration in the Ã^{′2}Δ state. Measured rotational lines have been assigned and rotational and fine structure parameters determined through a combined least-squares fit with the millimeter-wave pure rotational data of the X̃^{2}Σ^{+} state. Previous analyses of the Ã^{2}Π − X̃^{2}Σ^{+} transitions of BaOH and BaOD yielded significantly different spin-orbit coupling constants, which were attributed to possible global and local perturbations arising from vibrationally excited bands of the Ã^{′2}Δ state. Although the newly observed Ã^{′2}Δ state bands could not be conclusively designated a specific spin state, the derived Λ-doubling constants also show significant ^{2}Π character, further indicating a strong interaction between the Ã^{2}Π and Ã^{′2}Δ states of BaOH. To validate these conclusions, ab initio calculations have been carried out to further understand the nature of the BaOH excited states. The wavefunctions of the D̃^{′2}Σ^{+}, D̃^{2}Σ^{+}, C̃^{2}Π, B̃^{2}Σ^{+}, Ã^{2}Π, Ã^{′2}Δ and X̃^{2}Σ^{+} states have been optimised with a state averaged multiconfigurational calculation using the MolPro software. Calculated vertical term energies show relatively good agreement with existing optical data.